36.5 mL of 0.18 mol/L NaOH are required to titrated 10 mL of nitric acid. Determine the concentration of nitric acid.
1
Expert's answer
2017-03-29T07:27:22-0400
NaOH + HNO3 = NaNO3 + H2O
C(NaOH) · V (NaOH) = C (NaNO3) · V (NaNO3)
35.5 · 0.18 = 10 · V (NaNO3)
V (NaNO3) = (35.5 · 0.18)/10 = 0.639 mol/L